OpenStudy (dollbeautyx):

A surveyor is estimating the distance across a river. The actual distance is 279.8m. The surveyor's estimate is 300m. Absolute error= ?m Percent error= ?%

jhonyy9 (jhonyy9):

abs.err. 300-279,8 = - i think or my opinion -perc. err. using the rule of 3 simple so what is your opinion about it ?

OpenStudy (dollbeautyx):

@hartnn

hartnn (hartnn):

'the absolute error is the magnitude of the difference between the exact value and the approximation.' ~Wiki. So it is just \(|300-279.8| = ..\)

OpenStudy (dollbeautyx):

what would the percentage be?

hartnn (hartnn):

%error = \(\dfrac{|approx \: value - exact \: value|}{|exact\: value|}\times 100 = \dfrac{absolute \: error}{|exact value|}\times 100\)

OpenStudy (dollbeautyx):

7.2?

hartnn (hartnn):

yes, 7.2 is correct :)

OpenStudy (dollbeautyx):

okay, you were a big help. Thanks :)

hartnn (hartnn):

welcome ^_^